OpenStudy (anonymous):

express 84 as a product of its prime factors

OpenStudy (anonymous):

Keep dividing by the smallest prime number that will go into what you have left, and you will get the factorization. So, in this case\[84=2(42)=2^2(21)=2^2*3*7\]